OpenStudy (anonymous):

the slope int form equation is y=1/2x+5/2 how do i put into standard (general) form

OpenStudy (anonymous):

Ax+By=C is standard form

OpenStudy (anonymous):

so will it be..

OpenStudy (anonymous):

1/2x + y = 5/2 ?

OpenStudy (anonymous):

or do i have to change signs around

OpenStudy (anonymous):

on the right track, I would do : y=1/2x+5/2 2y=x+5 -x+2y=5 You have to change the sign for your x because you have to subtract on both sides, just remember when you +-x/ on both sides of the = :) hope that helped

OpenStudy (anonymous):

to check to see if your answer is correct, you could always do the math and turn it back into y=, if you get the same answer that you started with than you got it right.

OpenStudy (anonymous):

howd you get 2y=x+5